OpenStudy (austinl):

Slopes of Tangent Lines of Polar Equations a) Find all points where the following curves have vertical and horizontal tangent lines. b) Find the slope of the lines tangent to the curve at the origin (when relevant) c) Sketch the curve and all the tangent lines identified in parts (a) and (b).2\theta \[r = 2\cos(2 \theta)\]

OpenStudy (abb0t):

slope of tangent means derivative. find it for \(\large \frac{dr}{d \theta}\)

OpenStudy (austinl):

I know how to find a slope.... I need to figure out on the graph where these points are.

OpenStudy (abb0t):

You could convert it to cartesian coordinates (x,y) to make it easier to graph.

OpenStudy (austinl):

 It looks similar to this, devilishly hard to graph well on here.

OpenStudy (fibonaccichick666):

maybe think about, when does the derivative =0? and how many radians before it repeats itself

OpenStudy (austinl):

That, that actually helped.

OpenStudy (fibonaccichick666):

happy it did :)

OpenStudy (austinl):

Ok, so I have... Vertical: (2,pi/2),(2,3pi/2) Horizontal: (2,0),(2,pi) This correct?

OpenStudy (fibonaccichick666):

What about for \((\infty, \infty)\)?

OpenStudy (fibonaccichick666):

sorry (-∞,∞)